# Radiated energy fraction of SPI-induced disruptions at ASDEX Upgrade

### v1

@@ 3.3 Additional heating power $\text{W}_{\text{heat}}$
The term $\text{W}_{\text{heat}}$ is added to the calculation as the heating systems typically do not shut down instantly at the time of the FL, but continue heating the plasma (compare plasma heating after $\text{t}_{\text{FL}}$ in figure 4(b)). For this calculation it is assumed, that all the energy of the heat sources is absorbed. Automatic shut-down of external heating systems typically occurs when each system detects unsatisfactory absorption (i.e. NBI shine-through or ECRH reflection). Therefore, $\text{W}_{\text{heat}}$ is calculated as
$\text{W}_{\text{heat}}=\int_{\text{t}_{\text{FL}}}^{\text{t}_{\text{end}}}P_{\text{NBI}}+P_{\text{ECRH}}+P_{\text{Ohmic}},$ (6)
@@
$P_{\text{Ohmic}}=I_{\text{P}}\cdot U_{\text{loop}}.$ (7)
The $U_{\text{loop}}$ measurement at the plasma edge is meant to be representative of the Ohmic power input from the central solenoid during the disruption. On the time scales considered, the electric field induced in the core typically does not propagate to the edge measurement [40, 41].
@@ 3.4 Coupled energy $\text{W}_{\text{c}}$
As there is no direct measurement of the entire coupled energy $\text{W}_{\text{c}}$ into the surrounding structure, $\text{W}_{\text{c}}$ is estimated as 50% of the pre-injection magnetic energy, based on the observations by Sheikh et al [36], Lehnen et al [42], and JOREK simulations for ASDEX Upgrade VDEs performed by Schwarz [8] shown in figure 3. In order to estimate the coupled energy, 2D simulations with the extended MHD code JOREK [43] were carried out based on an L-mode discharge [7]. Different quantities of neon were introduced with a uniform source in a post-TQ plasma, leading to different CQ durations. One can calculate the radiated energy and the dissipated energy by Ohmic heating $\text{W}_{\text{Ohmic}}$ during the CQ phase. As the magnetic energy can either be lost by Ohmic heating or by coupling to the magnetic structures, the coupled energy is estimated as:
$\text{W}_{\text{c}}=\text{W}_{\text{mag}}-\text{W}_{\text{Ohmic}}.$ (8)
@@
The results are shown in figure 3, where the fraction of the magnetic energy that is coupled to the EM structures ($\text{f}_{\text{c}}$) and the radiation fraction of the available magnetic energy ($\text{f}_{\text{rad}}$) against $\Delta\text{t}_{CQ}$ are shown. Note, that in the simulation the external heating sources (NBI, ECRH) are stopped at the start of the disruption. As the neon source is added after the TQ, these quantities only refer to the CQ phase. These simulations can be used as a basis to validate a lumped parameter model of the conductive structures to estimate the coupled energy as done by Lehnen et al [42]. For low neon content ($1-5\times 10^{18}$ atoms) the disruptions are no longer radiation dominated, a direct comparison with the experimental results becomes increasingly difficult, and beyond the scope of this simple $\text{W}_{\text{c}}$ estimate. The duration of the CQ is calculated between the IP-spike (100%) and the end of the CQ (0%) [37] for the experiments and simulations for better comparability. A wall time of about 60 ms is used in these simulations [8]. We expect the coupled energy $\text{W}_{\text{c}}$ to depend strongly on the wall time for the net vessel current. The shorter this time is, the less energy is coupled into the vessel. At the time these simulations were performed, the complicated wall structures were not well represented in the coupled JOREK-STARWALL simulations. Using the toroidal resistivity of $3\ \mu\Omega\textrm{m}$ derived by Giannone et al [44] instead of the $0.45\ \mu\Omega\textrm{m}$ used in the simulations, we expect the wall time to be in the range of $\sim 9$ ms instead. While for short CQ durations ($<10$ ms) the 50% coupled energy remains a good approximation, the coupled energy may be overestimated for longer CQ durations as observed for the low neon and pure deuterium SPI. Ultimately, the radiated energy fraction of these discharges might be lower.
@@ 3.5 Plasma scenarios of the 2022 SPI experiments
Figure 4: Overview plot for the “SPI H-mode” scenario from table 3. In (a) the entire discharge is provided, while in (b) the injection and disruption part is shown in more detail. Note the different y-axis limits between (a) and (b) in (II) and (VI). The following signals are provided: In (I), plasma current, stored magnetic and thermal energies are provided. The line integrated ($\textrm{m}^{-2}$) electron densities derived from the core DCN [45, 46] and fast CO2 lasers [47] are displayed in (II). Hereby, the fast CO2 laser signals are smoothed via a rolling window average with window size of $\textrm{w}_{\textrm{avg}}=16$ ms in (a). In (III), the heating powers (NBI, ECRH, Ohmic) are shown. In (IV), the electron temperature via the electron cyclotron emission (ECE) [48] measurement for different radial channels, with $\textrm{w}_{\textrm{avg}}=3.2$ ms in (a) as well as the $\textrm{T}_{\textrm{e}}$ reconstruction from the integrated data analysis (IDA) [49], with $\textrm{w}_{\textrm{avg}}=5$ ms in (a). Note, that the ECE signals go into cut-off at $\approx 2.315$ sec, hence do not reflect the TQ time (compare $\text{W}_{\text{th}}$ in (I)). In (V), the core and edge (at $\rho_{\textrm{pol}}=0.95$) safety factors are given. In (VI), the total radiated power in the five toroidal sectors derived from the foil bolometers (equation (1)) is provided, with $\textrm{w}_{\textrm{avg}}=20$ ms in (a). In (VIb), the total radiated energy (equation (3)) and radiated energy fraction (equation (2)) are provided. In this case, three large radiation peaks are visible at: fragment arrival, the TQ/IP-spike, and the final VDE phase.
@@
SPI H-mode 1.8 7.5–9.9 2.85–3.2 3.74–3.87 4.5–6.8 1.25–1.4 0.48
2.5T H-mode 2.5 9.3–9.9 2.85–3.0 5.24–5.26 5.5–6.8 1.3–1.4 0.48
Low Te 1.8 2.4 1.6 3.78–3.80 5.5–6.1 0.93 0.33
High $\text{W}_{\text{plasma}}$ 2.5 8.3–9.7 2.05 4.09–4.12 no measurement 1.92 0.47
Ohmic 1.8 3.55–3.68 3.9–4.5 0.69–0.74 0.15
@@ 4 Experimental values of $\text{W}_{\text{rad}}$ and $\text{f}_{\text{rad}}$
Figure 5: The radiated energy during SPI, plotted against the plasma energy (see equation (2)) at $\text{t}_{\text{FL}}$ and $\text{t}_{\text{end}}$, respectively. The five different target plasmas indicated in (a) are introduced in table 3. The colour of the points represents the amount of neon inside the pellets as indicated by the colourbar in the center. The estimated number of neon atoms (left) which sets the point colour and the equivalent partial pressure ratio in the mix tank for the 8 mm pellets (right). The number inside the bracket represents the number of pellets (including 4 and 8 mm pellets).
In this section, the $\text{W}_{\text{rad}}$ and resulting $\text{f}_{\text{rad}}$ values for the 2022 experimental campaign at ASDEX Upgrade are presented. Figure 5 shows the total radiated energy $\text{W}_{\text{rad}}$ plotted against the plasma stored energy $\text{W}_{\text{plasma}}$ as defined in equation (2). The colour-coding shows the neon content inside the pellet ranging from pure deuterium to pure neon pellets on a logarithmic scale, markers represent the 4 mm and 8 mm pellets for (non-)disruptive injections. Numbers in the brackets indicate the number of pellets in that category. Note, that while the neon partial pressure ratio (right side of the colourbar) only corresponds to the 8 mm pellets, the number in the brackets indicate the total number of 4 and 8 mm pellets with this neon percentage. The target plasma scenarios from section 3.5 are indicated in figure 5(a).
In figure 5(a) the plasma stored energy is evaluated for the time of the first light, hence without the additional $\text{W}_{\text{heat}}$ term, consequently the scenarios occur at distinct values of $\text{W}_{\text{plasma}}$. The H-mode scenarios at 1.8 T (SPI H-mode) and 2.5 T share the same $\text{W}_{\text{plasma}}$. As expected, increasing the amount of neon inside the pellet leads to an increase in radiated energy fraction from around 10% (4 mm diameter pure D2 pellets) to 100% (high neon concentrations of 10% and beyond for the 8 mm diameter pellets). Without additional heating, the total plasma stored energy is given by $\text{W}_{\text{plasma}}$. The total available energy including external heating is summarised as the term $\text{W}_{\text{plasma}}$ ($\text{t}_{\text{end}}$) (figure 5(b)). Hereby, long disruption phases (lower neon content) are shifted further to the right, as typically more heat is injected into the plasma until the heating systems are switched off. The radiated energy fraction of non-disruptive SPI is below 30% (indicated by the blue area in figure 5).
@@
With increasing neon content inside the pellet, the radiated energy fraction also increases. However, $\text{f}_{\text{rad}}$ saturates around 80% when we take the $\text{W}_{\text{heat}}$ into account. Whether this energy is not absorbed by the plasma in the first place, conducted to the PFCs as heat, or if potential other loss mechanisms are at play, is presently not identified. Note, that the coupled energy $\text{W}_{\text{c}}$ for the low neon content/pure deuterium with long CQ times might be overestimated from the present simulations (see section 3.4), which would result in lower radiated energy fractions for these injections. For disruptive SPI, the radiated energy fraction is between 30% (pure deuterium SPI) and 80–85% ($>10^{21}$ neon atoms). This is in line with previous results from JET SPI, where most radiated energy fractions were between 30% and 85% [51]. However, assessing the radiated energy fraction in JET is challenging given the bolometric setup, hence either toroidally symmetric radiation is assumed [51], different toroidal mode-locking positions are examined [52] or 3D simulation codes like Emis3D are applied [51, 53].
@@
For two cases, a disruption was triggered by pure deuterium massive gas injection (MGI) and an 8 mm SPI pellet with 10% neon was injected into the disrupting plasma (purple triangles). In the first realisation, the fragments of the SPI arrived after the thermal collapse during the thermal quench (TQ), hence the radiated energy (and radiated energy fraction) is low. With reduced time delay between MGI and SPI, a higher $\text{W}_{\text{rad}}$ ($\text{f}_{\text{rad}}$) value was achieved.
@@
The magenta points in figure 6 indicate discharges in which a discharge control system (DCS) [54, 55, 56] action was triggered by the SPI (DCS/VPEState changed from state 0 to a different state) during the pre-TQ phase. If the DCS was unable to recover the discharge [54] and detects a plasma state that could lead to a disruption (e.g. the plasma current center position passes a $z_{\textrm{current}}$-position threshold to detect VDEs) it goes into a “holding” state (indicated by the state-change of the VPEState), trying to ramp down the discharge safely if the plasma does not disrupt before the DCS can take any action. Hereby, in most cases the threshold of the allowed vertical displacement of the plasma was exceeded which triggered the state-change of the DCS. The resulting action of the control system might have affected the development of some of these disruptions and in some cases we understand, that the attempt to ramp down the plasma even caused the disruption for injections which – if the threshold value of the VPEState would have been set to higher values – might be recoverable via the DCS: We observe that some discharges (either due to interlocks in the NBI or the VPEState-change and consequent holding-action) do disrupt which otherwise were already about to recover where the $z_{\textrm{current}}$ position was slowly going back to pre-injection levels and below the trigger value of the VPEState. One of the implications of the holding-state is, that only three NBI sources are allowed to soft-land the discharge (ramp-down), while four were requested for our scenario, i.e. in these cases one NBI source is switched off. However, without the desired beam power for the specific scenario during the holding-state, this might be actually triggering the disruption. Eventually, at some point inside the CQ the VPEState-change may be triggered, however, only discharges where the VPEState changed before the IP-spike are selected for the analysis in this work and excluded from figures 357, and 8. However, it was observed in some cases (potentially all/a large fraction of cases are affected) that all NBI sources are switched-off shortly before the VPEState-change of the DCS is triggered due to interlocks in the NBI system (e.g. to prevent excessive shine-through). As a consequence a more detailed study on the impact of the switch-off of the heating systems on the pre-TQ time would be required.
@@
Figure 7: (a) Radiated energy as a function of time and (b) estimated number of deuterium molecules ($\text{D}_{2}$) for pure deuterium SPI into the SPI H-mode. Only discharges are displayed, in which the VPEState-change of the DCS happened at or after the IP-spike. In (a), a subset of the H-mode injections is shown with the number of deuterium molecules as the colour-code. The time $t=0$ ms is with respect to the start of the integration interval (before the first radiation peak) which is around $\text{t}_{\text{FL}}$. The radiation increase for $t<0$ indicates the pre-injection radiation levels, hence is similar to the radiated power after the recovery for non-disruptive SPI (compare slopes). Three branches emerge with the (full length) 4 mm, 8 mm shortened pellets and the full length 8 mm pellet injections. The threshold of disruptive injections is around $1\times 10^{22}$ deuterium molecules in line with previous simulation results [50]. The discharge #40657 exhibits a long pre-TQ phase and might be close to the threshold of being non-disruptive.
@@
The pure deuterium, 4 mm diameter (or even a few low neon doped) pellets did typically not cause the SPI H-mode plasmas to disrupt. While the full-sized, 8 mm, pure deuterium pellets caused a disruption, the shortened (approx. half nominal pellet length [24]) pellets did not. For an excerpt of pure deuterium SPI in figure 7, the time evolution of the radiated energy with the colour-coding representing the estimated number of deuterium molecules is shown in (a). In figure 7(b) the radiated energy at the end of the integration interval is plotted as a function of estimated number of deuterium molecules. Note, that for some pellets (especially the 4 mm) no measurement of the pellet length was available via the OPD camera. For these pellets, the estimated number of deuterium molecules is taken from pellets with the same recipe for desublimation and launching from the AUG SPI database [24] which results in the same number of estimated deuterium molecules (vertically aligned points). Estimation from the gas reservoir pressure drop alone is difficult due to the various heatings applied [24]. Only discharges which had their VPEState change of the DCS after the pre-TQ phase (during the IP-spike time window or CQ phase) are included in figure 7. For the discharge #40657 ($25^{\circ}$, circular shatter head; $\sim 460$ m/s pre-shatter velocity; see SPI H-mode overview figure 4) we do not have a direct measurement of the pellet length, hence the value of around $1.4\times 10^{21}$ deuterium molecules is inferred from the SPI recipe database. However, this discharge exhibits a long pre-TQ phase (around 10 ms) compared to the other injections depicted in figure 7(a). Consequently, we suspect this discharge to be close to being non-disruptive. The VPEState change happened for this discharge right at the IP-spike (which is very early compared to other discharges displayed in figure 7 where at some point during the CQ the VPEState-change is triggered), hence the discharge would probably be terminated by the loss of one of the NBI sources due to the control system otherwise. The observation of a disruption threshold around $1\times 10^{21}$ deuterium molecules is well in line with simulation results by Hölzl et al [50]. Scanning the amount of deuterium inside the pellets, the simulation results suggest incomplete TQs below a threshold of $(8\pm 4)\times 10^{21}$ deuterium atoms. Note, that the difference of an factor of two between the simulations and experiments might be also causethd by the simulation setup: In the simulations, all of the material was assumed to be injected in the form of solid fragments, while in the experiment, part of the pellet material will be transformed into gas during the shattering process which might not (quickly) penetrate deeply into the plasma. Compared to the threshold of the experiments for H-modes, we observed Ohmic discharges to disrupt already for the 4 mm diameter pellets (see figure 7(b)).

### v2

@@ 3.3 Additional heating power $\text{W}_{\text{heat}}$
The term $\text{W}_{\text{heat}}$ is added to the calculation as the heating systems typically do not shut down instantly at the time of the FL, but continue heating the plasma (compare plasma heating after $\text{t}_{\text{FL}}$ in figure 5(b)). For this calculation it is assumed, that all the energy of the heat sources is absorbed. Due to this assumption, the resulting values of $\text{f}_{\text{rad}}$ in figure 6(b) can be interpreted as a lower limit compared to the definition of $\text{W}_{\text{plasma}}$ without $\text{W}_{\text{heat}}$, which may serve as an upper limit. Automatic shut-down of external heating systems typically occurs when each system detects unsatisfactory absorption (i.e. NBI shine-through or ECRH reflection). Therefore, $\text{W}_{\text{heat}}$ is calculated as
$\text{W}_{\text{heat}}=\int_{\text{t}_{\text{FL}}}^{\text{t}_{\text{end}}}P_{\text{NBI}}+P_{\text{ECRH}}+P_{\text{external Ohmic}},$ (6)
@@
$P_{\text{external Ohmic}}=I_{\text{P}}\cdot U_{\text{loop}}.$ (7)
The $U_{\text{loop}}$ measurement at the plasma edge is meant to be representative of the external Ohmic power input from the central solenoid during the disruption. On the time scales considered, the electric field induced in the core typically does not propagate to the edge measurement [43, 44].
@@ 3.4 Coupled energy $\text{W}_{\text{c}}$
As there is no direct measurement of the entire coupled energy $\text{W}_{\text{c}}$ into the surrounding structure, $\text{W}_{\text{c}}$ is estimated as 50% of the pre-injection magnetic energy, based on the observations by Sheikh et al [38] and JOREK simulations for ASDEX Upgrade VDEs performed by Schwarz [8] shown in figure 4. (We note, that, coincidentally, a value of 50% was also estimated for JET by Lehnen et al [37].) In order to estimate the coupled energy, 2D simulations with the extended MHD code JOREK [45] were carried out based on an L-mode discharge [7]. Different quantities of neon were introduced with a uniform source in a post-TQ plasma, leading to different CQ durations. One can calculate the radiated energy and the dissipated energy by Ohmic heating $\text{W}_{\text{internal Ohmic}}$ during the CQ phase. As the magnetic energy can either be lost by internal Ohmic heating or by coupling to the magnetic structures (and the RE population is negligible), the coupled energy is estimated as:
$\text{W}_{\text{c}}=\text{W}_{\text{mag}}-\text{W}_{\text{internal Ohmic}}.$ (8)
@@
The results are shown in figure 4, where the fraction of the magnetic energy that is coupled to the conducting structures ($\text{f}_{\text{c}}$) and the radiation fraction of the available magnetic energy ($\text{f}_{\text{rad}}$) against $\Delta\text{t}_{CQ}$ are shown. Note, that in the simulation the external heating sources (NBI, ECRH) are stopped at the start of the disruption. As the neon source is added after the TQ, these quantities only refer to the CQ phase. These simulations can be used as a basis to validate a lumped parameter model of the conductive structures to estimate the coupled energy as done by Lehnen et al [37]. For low neon content ($1-5\times 10^{18}$ atoms) the disruptions are no longer radiation dominated, a direct comparison with the experimental results becomes increasingly difficult, and beyond the scope of this simple $\text{W}_{\text{c}}$ estimate. The duration of the CQ is calculated between the IP-spike (100%) and the end of the CQ (0%) [39] for the experiments and simulations for better comparability. A wall time of about 60 ms is used in these simulations [8]. We expect the coupled energy $\text{W}_{\text{c}}$ to depend strongly on the wall time for the net vessel current. The shorter this time is, the less energy is coupled into the vessel. At the time these simulations were performed, the complicated wall structures were not well represented in the coupled JOREK-STARWALL simulations. Using the toroidal resistivity of $3\ \mu\Omega\textrm{m}$ derived by Giannone et al [46] instead of the $0.45\ \mu\Omega\textrm{m}$ used in the simulations, we expect the wall time to be in the range of $\sim 9$ ms instead. While for short CQ durations ($<10$ ms) the 50% coupled energy remains a good approximation, the coupled energy may be overestimated for longer CQ durations as observed for the low neon and 100% $\textrm{D}_{2}$ SPI. Ultimately, the radiated energy fraction of these discharges from section 4 might be lower.
@@ 3.5 Plasma scenarios of the 2022 SPI experiments
Figure 5: Overview plot for the “SPI H-mode” scenario from table 3. In (a) the entire discharge is provided, while in (b) the injection and disruption part is shown in more detail. The disruption is initiated by injection of a 100% $\textrm{D}_{2}$ injection, triggered at 2.30 s. Note the different y-axis limits between (a) and (b) in (II) and (VI). The following signals are provided: In (I), plasma current, stored magnetic and thermal energies are provided. The line integrated ($\textrm{m}^{-2}$) electron densities derived from the core DCN [47, 48] and fast CO2 lasers [49] are displayed in (II). Hereby, the fast CO2 laser signals are smoothed via a rolling window average with window size of $\textrm{w}_{\textrm{avg}}=16$ ms in (a). In (III), the heating powers (NBI, ECRH, Ohmic) are shown. In (IV), the electron temperature via the electron cyclotron emission (ECE) [50] measurement for different radial channels, with $\textrm{w}_{\textrm{avg}}=3.2$ ms in (a) as well as the $\textrm{T}_{\textrm{e}}$ reconstruction from the integrated data analysis (IDA) [51], with $\textrm{w}_{\textrm{avg}}=5$ ms in (a). Note, that the ECE signals go into cut-off at $\approx 2.315$ sec, hence do not reflect the TQ time (compare $\text{W}_{\text{th}}$ in (I)). In (V), the core and edge (at $\rho_{\textrm{pol}}=0.95$) safety factors are given. In (VI), the total radiated power in the five toroidal sectors derived from the foil bolometers (equation (1)) is provided, with $\textrm{w}_{\textrm{avg}}=20$ ms in (a). In (VIb), the total radiated energy (equation (3)) and radiated energy fraction (equation (2)) are provided. In this case, three large radiation peaks are visible at: fragment arrival, the TQ/IP-spike, and the final VDE phase.
@@
SPI H-mode 1.8 7.5–9.9 2.85–3.2 $<4$ 4.5–6.8 1.25–1.4 0.48
2.5T H-mode 2.5 9.3–9.9 2.85–3.0 $>5$ 5.5–6.8 1.3–1.4 0.48
Low Te 1.8 2.4 1.6 $<4$ 5.5–6.1 0.93 0.33
High $\text{W}_{\text{plasma}}$ 2.5 8.3–9.7 2.05 $\approx 4$ no measurement 1.92 0.47
Ohmic 1.8 $<4$ 3.9–4.5 0.69–0.74 0.15
@@ 4 Experimental values of $\text{W}_{\text{rad}}$ and $\text{f}_{\text{rad}}$
Figure 6: The radiated energy during SPI, plotted against the energy dissipated in the plasma $\text{W}_{\text{plasma}}$ [37] (see equation (2)) at $\text{t}_{\text{FL}}$ and $\text{t}_{\text{end}}$, respectively. The five different target plasmas indicated in (a) are introduced in table 3. The colour of the points represents the amount of neon inside the pellets as indicated by the colourbar in the center. The estimated number of neon atoms (left colour-scale) which sets the point colour and the equivalent partial pressure ratio in the mix tank for the 8 mm pellets (right). The number inside the bracket represents the number of pellets (including 4 and 8 mm pellets). The $\text{W}_{\text{heat}}$-term in (b) shifts $\text{W}_{\text{plasma}}$ to higher values (to the right) indicated by the arrows and red circled value, depending on pre-TQ duration. Therefore, an upper limit on the $\text{f}_{\text{rad}}$-values is provided in (a) and a lower limit in (b) as this assumes 100% absorbtion of $\text{W}_{\text{heat}}$.
In this section, the $\text{W}_{\text{rad}}$ and resulting $\text{f}_{\text{rad}}$ values for the 2022 experimental campaign at ASDEX Upgrade are presented. Figure 6 shows the total radiated energy $\text{W}_{\text{rad}}$ plotted against the energy dissipated in the plasma $\text{W}_{\text{plasma}}$ as defined in equation (2). The colour-coding shows the neon content inside the pellet ranging from 100% $\textrm{D}_{2}$ to 100% Ne pellets on a logarithmic scale, markers represent the 4 mm and 8 mm pellets for (non-)disruptive injections. Numbers in the brackets indicate the number of pellets in that category. Note, that while the neon partial pressure ratio ($\textrm{f}_{\textrm{neon}}$ – right side of the colourbar) only corresponds to the 8 mm pellets, the number in the brackets indicate the total number of 4 and 8 mm pellets with this neon percentage. This partial pressure ratio ($\textrm{f}_{\textrm{neon}}$) for the 8 mm pellets is used for the 2nd x-axis in figures 7 and 9. The target plasma scenarios from section 3.5 are indicated in figure 6(a).
In figure 6(a) the plasma stored energy is evaluated for the time of the first light, hence without the additional $\text{W}_{\text{heat}}$ term, consequently the scenarios occur at distinct values of $\text{W}_{\text{plasma}}$. The H-mode scenarios at 1.8 T (SPI H-mode) and 2.5 T share the same $\text{W}_{\text{plasma}}$. As previously reported for DIII-D [53], increasing the amount of neon inside the pellet leads to an increase in radiated energy fraction from around 10% (4 mm diameter 100% $\textrm{D}_{2}$ pellets) to 100% (high neon concentrations of 10% and beyond for the 8 mm diameter pellets). Without additional heating, the energy dissipated in the plasma given by $\text{W}_{\text{plasma}}$ [37]. The total dissipated energy including external heating is summarised as the term $\text{W}_{\text{plasma}}$ ($\text{t}_{\text{end}}$) (figure 6(b)). Hereby, long disruption phases (lower neon content) are shifted further to the right, as typically more heat is injected into the plasma until the heating systems are switched off. The radiated energy fraction of non-disruptive SPI is below 30% (indicated by the blue area in figure 6).
@@
With increasing neon content inside the pellet, the radiated energy fraction also increases. However, $\text{f}_{\text{rad}}$ saturates around 80% when we take the $\text{W}_{\text{heat}}$ into account. Whether this energy is not absorbed by the plasma in the first place, conducted to the PFCs as heat, or if potential other loss mechanisms are at play, is presently not identified. Note, that the coupled energy $\text{W}_{\text{c}}$ for the low neon content/100% $\textrm{D}_{2}$ with long CQ times might be overestimated from the present simulations (see section 3.4), hence the actual radiated energy fractions for these injections may be lower. For disruptive SPI, the radiated energy fraction is between 30% (100% $\textrm{D}_{2}$ SPI) and 80–85% ($>10^{21}$ neon atoms). This is in line with previous results from JET SPI, where most radiated energy fractions were between 30% and 85% [54]. However, assessing the radiated energy fraction in JET is challenging given the bolometric setup, hence either toroidally symmetric radiation is assumed [54], different toroidal mode-locking positions are examined [55] or 3D simulation codes like Emis3D are applied [54, 56]. For DIII-D, radiated fractions of $\text{W}_{\text{rad}}/\text{W}_{\text{th}}\lessapprox 90\%$ are reported [53].
@@
For two cases, a disruption was triggered by 100% $\textrm{D}_{2}$ massive gas injection (MGI) and an 8 mm SPI pellet with 10% neon was injected into the disrupting plasma (purple triangles). In the first realisation, the fragments of the SPI arrived after the thermal collapse during the thermal quench (TQ), hence the radiated energy (and radiated energy fraction) is low. With reduced time delay between MGI and SPI, a higher $\text{W}_{\text{rad}}$ ($\text{f}_{\text{rad}}$) value was achieved.
@@
The magenta points in figure 7 indicate discharges in which a discharge control system (DCS) [57, 58, 59] action was triggered by the SPI (DCS/VPEState changed from state 0 to a different state) during the pre-TQ phase. If the DCS was unable to recover the discharge [57] and detects a plasma state that could lead to a disruption (e.g. the plasma current center position passes a $z_{\textrm{current}}$-position threshold to detect VDEs) it goes into a “holding” state (indicated by the state-change of the VPEState), before trying to ramp down the discharge safely if the plasma does not disrupt before the DCS can take any action. Hereby, in most cases the threshold of the allowed vertical displacement of the plasma was exceeded which triggered the state-change of the DCS. The resulting action of the control system might have affected the development of some of these disruptions and in some cases we understand, that the attempt to ramp down the plasma even caused the disruption for injections which – if the threshold value of the VPEState would have been set to higher values – might be recoverable via the DCS: We observe that some discharges (either due to interlocks in the NBI or the VPEState-change and consequent holding-action) do disrupt which otherwise were already about to recover where the $z_{\textrm{current}}$ position was slowly going back to pre-injection levels and below the trigger value of the VPEState. One of the implications of the holding-state is, that only three NBI sources are allowed to soft-land the discharge (ramp-down), while four were requested for our scenario, i.e. in these cases one NBI source is switched off. However, without the desired beam power for the specific scenario during the holding-state, this might be actually triggering the disruption. Eventually, at some point inside the CQ the VPEState-change may be triggered, however, only discharges where the VPEState changed before the IP-spike are selected for the analysis in this work and excluded from figures 468, and 9. However, it was observed in some cases (potentially all/a large fraction of cases are affected) that all NBI sources are switched-off shortly before the VPEState-change of the DCS is triggered due to interlocks in the NBI system (e.g. to prevent excessive shine-through). As a consequence a more detailed study on the impact of the switch-off of the heating systems on the pre-TQ time would be required.
@@
Figure 8: (a) Radiated energy as a function of time and (b) estimated number of deuterium molecules ($\text{D}_{2}$) for 100% $\textrm{D}_{2}$ SPI into the SPI H-mode. Only discharges are displayed, in which the VPEState-change of the DCS happened at or after the IP-spike. In (a), a subset of the H-mode injections is shown with the number of deuterium molecules as the colour-code. The time $t=0$ ms is with respect to the start of the integration interval (before the first radiation peak) which is around $\text{t}_{\text{FL}}$. The radiation increase for $t<0$ indicates the pre-injection radiation levels, hence is similar to the radiated power after the recovery for non-disruptive SPI (compare slopes). Three branches emerge with the (full length) 4 mm, 8 mm shortened pellets and the full length 8 mm pellet injections. The threshold of disruptive injections is around $1\times 10^{22}$ deuterium molecules in line with previous simulation results [52]. Note, this disruption threshold also depends strongly on the actual material assimilation, hence the fragmentation & ablation properties for this discharge, with the number of injected particles used as an approximation. The discharge #40657 exhibits a long pre-TQ phase – compare other disruptive cases in red-orange in (a) – and might be close to the threshold of being non-disruptive (yellow-green lines).
@@
The 100% $\textrm{D}_{2}$, 4 mm diameter (or even a few low neon doped) pellets did typically not cause the SPI H-mode plasmas to disrupt. While the full-sized, 8 mm, 100% $\textrm{D}_{2}$ pellets caused a disruption, the shortened (approx. half nominal pellet length [24]) pellets did not. For an excerpt of 100% $\textrm{D}_{2}$ SPI in figure 8, the time evolution of the radiated energy with the colour-coding representing the estimated number of deuterium molecules is shown in (a). In figure 8(b) the radiated energy at the end of the integration interval is plotted as a function of estimated number of deuterium molecules. Note, that for some pellets (especially the 4 mm) no measurement of the pellet length was available via the OPD camera. For these pellets, the estimated number of deuterium molecules is taken from pellets with the same recipe for desublimation and launching from the AUG SPI database [24] which results in the same number of estimated deuterium molecules (vertically aligned points). Estimation from the gas reservoir pressure drop alone is difficult due to the various heatings applied [24]. Only discharges which had their VPEState change of the DCS after the pre-TQ phase (during the IP-spike time window or CQ phase) are included in figure 8. For the discharge #40657 ($25^{\circ}$, circular shatter head; $\sim 460$ m/s pre-shatter velocity; see SPI H-mode overview figure 5) we do not have a direct measurement of the pellet length, hence the value of around $1.4\times 10^{21}$ deuterium molecules is inferred from the SPI recipe database. However, this discharge exhibits a long pre-TQ phase (around 10 ms) and also lower initial radiation during the pre-TQ phase compared to the other injections depicted in figure 8(a). Consequently, we suspect this discharge to be close to being non-disruptive. The VPEState change happened for this discharge right at the IP-spike (which is very early compared to other discharges displayed in figure 8 where at some point during the CQ the VPEState-change is triggered), hence the discharge would probably be terminated by the loss of one of the NBI sources due to the control system otherwise. The observation of a disruption threshold around $1\times 10^{22}$ deuterium molecules is well in line with simulation results by Hölzl et al [52]. Scanning the amount of deuterium inside the pellets, the simulation results suggest incomplete TQs below a threshold of $(8\pm 4)\times 10^{21}$ deuterium atoms. Note, that the difference of an factor of two between the simulations and experiments might be also caused by the simulation setup: In the simulations, all of the material was assumed to be injected in the form of solid fragments, while in the experiment, part of the pellet material will be transformed into gas during the shattering process which might not (quickly) penetrate deeply into the plasma. Also the fragment size and velocity distribution will impact the amount of assimilated $\textrm{D}_{2}$, hence the disruption threshold. It depends on the assimilated pellet material, however, only indicated by the injected amount in figure 8(b). Compared to the threshold of the experiments for H-modes, we observed Ohmic discharges to disrupt already for the 4 mm diameter pellets (see figure 8(b)).